OpenStudy (anonymous):

[9.02] Where does the graph of y = 2x2 − x − 3 cross the x-axis?

OpenStudy (lgbasallote):

that would be the x-intercepts so let y = 0 then solve for x 0 = 2x^2 - x - 3 2x ^ 2- x - 3 = 0 (2x-3)(x+1) = 0 now solve for x to get your answer

OpenStudy (anonymous):

Thats where I'm getting confused. solving this for x.

OpenStudy (lgbasallote):

hint..when you have (x+2)(x-1) = 0 the roots will just be x + 2 = 0 and x - 1 =0 solve for x in each case when x + 2 = 0 x = -2 when x - 1 = 0 x = 1 does that help?

OpenStudy (anonymous):

yes thank you

OpenStudy (lgbasallote):

welcome ^_^